The amount A (t), in grams, of radioactive material in a sample after years, is given

by A^t=80(2^(-t/100)

(a) Find the amount of material in the original sample /1mrks

(b) calculate the half-life of the material. (The half-life is the time taken for half of

the original material to decay).


1
Expert's answer
2021-10-04T18:17:22-0400

(a) "A(0)=80*2^{-\frac{0}{100}}=80."


(b) "40=80*2^{-\frac{t}{100}}"

"2^{-1}=2^{-\frac{t}{100}}"

"\frac{t}{100}=1"

"t=100"

Half-life is 100 years